Jose is 3 years less than 4 times as old as his daughter. If the sum of their ages is 37, how old is Jose?
Answer by ankor@dixie-net.com(12695)

J = Joe's age; D = daughters age
:
Write an equation just like it says:
"Jose is 3 years less than 4 times as old as his daughter."
J + 3 = 4*D
:
"If the sum of their ages is 37,"
J + D = 37
D = (37 - J)
:
how old is Jose?
Substitute (37-J) for D in the 1st equation:
J + 3 = 4(37-J)
J + 3 = 148 - 4J
J + 4J = 148 - 3
5J = 145
J = 145/5
J = 29 yrs
:
D's age = 37 - 29 = 8 yrs
:
:
Check using the statement:
""Jose is 3 years less than 4 times as old as his daughter."
29 + 3 = 4*8